Solve- (3x+y-z)p + (x+y+z)q = 2(z+y)
1
Expert's answer
2020-06-29T19:15:08-0400

Before solving the problem, please note the problem has been changed as "(3x+y+z)p + (x+y+z)q = 2(z+y)" instead of "(3x+y-z)p + (x+y+z)q = 2(z+y)".


Consider, "(3x+y+z)p + (x+y+z)q = 2(z+y)".

The auxiliary equation is,

"\dfrac{dx}{3x+y+z}=\dfrac{dy}{x+y+z}=\dfrac{dz}{2(z+y)}."


Using the multipliers 1, -3, 1 each of the above ratio is equal to

"\dfrac{dx-3dy+dz}{3x+y+z-3x-3y-3z+2z+2y}=\dfrac{dx-3dy+dz}{0}"


Thus,

"dx-3dy+dz = 0"

Integrating we get, "x-3y+z=c_{1}."


Each of the ratio in the auxiliary equation is equal to

"\dfrac{dx+dy+dz}{3x+y+z+x+y+z+2z+2y}=\dfrac{dx-dy-dz}{3x+y+z-x-y-z-2z-2y}\\~\\
\dfrac{dx+dy+dz}{4(x+y+z)}=\dfrac{dx-dy-dz}{2(x-y-z)}\\~\\
\dfrac{dx+dy+dz}{2(x+y+z)}=\dfrac{dx-dy-dz}{(x-y-z)}\\~\\
\text{Integrating,} \\
\dfrac{1}{2}\ln(x+y+z)=\ln(x-y-z)+\ln c_{2}\\~\\
\ln\bigg(\dfrac{\sqrt{x+y+z}}{x-y-z}\bigg)=\ln c_{2}\\~\\
\dfrac{\sqrt{x+y+z}}{x-y-z} = c_{2}."


Hence the general solution is, "\phi(c_{1},c_{2})=0"

"\phi\bigg(x-3y+z, \dfrac{\sqrt{x+y+z}}{x-y-z}\bigg)=0"